### v2

# CASE – Condition-Aware Sentence Embeddings for Conditional Semantic Textual Similarity Measurement

Abstract
The meaning conveyed by a sentence often depends on the context in which it appears. Despite the progress of sentence embedding methods, it remains unclear as how to best modify a sentence embedding conditioned on its context. To address this problem, we propose Condition-Aware Sentence Embeddings (CASE), an efficient and accurate method to create an embedding for a sentence under a given condition. First, CASE creates an embedding for the condition using an Large Language Model (LLM), where the sentence influences the attention scores computed for the tokens in the condition during pooling. Next, a supervised non-linear projection is learnt to reduce the dimensionality of the LLM-based text embeddings. We show that CASE significantly outperforms previously proposed Conditional Semantic Textual Similarity (C-STS) methods on an existing standard benchmark dataset. We find that subtracting the condition embedding will consistently improve the C-STS performance of LLM-based text embeddings. Moreover, we propose a supervised dimensionality reduction method that not only reduces the dimensionality of the LLM-based embeddings, but also significantly improves their performance.
CASE – Condition-Aware Sentence Embeddings for
Conditional Semantic Textual Similarity Measurement
Gaifan Zhang1  Yi Zhou2  Danushka Bollegala1 1 University of Liverpool  2 Cardiff University sggzhan8@liverpool.ac.uk, zhouy131@cardiff.ac.uk, danushka@liverpool.ac.uk
§ 1 Introduction
Representing the meaning of a given sentence using a vector embedding is a fundamental task required by many Natural Language Processing (NLP) applications Conneau et al. (2017); Reimers and Gurevych (2019); Gao et al. (2021); Xu et al. (2023); Chen et al. (2023). Sentence embeddings are used to measure the Semantic Textual Similarity (STS) between two sentences Agirre et al. (2012, 2015, 2016).
Despite its importance, measuring STS between two sentences is a non-trivial task for humans, which is conditioned on what is being compared between the two sentences. For example, for the two sentences shown in Figure 1, human annotators would assign different similarity ratings, depending on what they are asked to compare (i.e., given conditions 1 and 2). Existing STS benchmarks do not specify the conditions under which two sentences must be compared for their semantic similarity. To address this limitation, Deshpande et al. (2023) proposed the C-STS task and a dataset, where the similarity between two sentences $s_{1}$ and $s_{2}$ is measured under two different conditions $c_{\rm low}$ and $c_{\rm high}$, which focus on different aspects of semantics in the two sentences, resulting in different similarity ratings. Many real-world applications can be seen as C-STS tasks such as ranking a set of documents retrieved for the same query in Information Retrieval (IRManning et al. (2008), comparing two answers for the same question in Question Answering (QARisch et al. (2021), or measuring the strength of a semantic relation between two entities in Knowledge Graph Completion (KGCYoo et al. (2024); Lin et al. (2024).
Figure 1: The two conditions focus on different information described in the two sentences. Human annotators rate the two sentences 1–5, indicating a high-level (5) of semantic textual similarity under condition 2 than condition 1 (1). Our proposed condition-aware sentence embedding (CASE) method reports similarity scores that are well-aligned with the human similarity ratings.
We propose CASE, a method for learning Condition-Aware Sentence Embeddings, for a given input sentence considering another sentence. Specifically, CASE creates an embedding for the condition sentence using an LLM with a prompt that includes the target sentence (e.g. Sentences 1/2 in Figure 1), where the latter is not encoded explicitly in the embedding but influences the attention scores used during token pooling. Compared to Masked Language Models that have been used extensively in prior work on C-STS, LLMs contain billions of parameters and are typically trained on much larger datasets for a longer period of time. Consequently, LLM-based embedding models have consistently ranked at the top in leader boards evaluating text embedding models Muennighoff et al. (2022). Therefore, by leveraging LLM-based embeddings, CASE is able to benefit from the rich world knowledge contained in LLMs. However, the optimal method to use LLMs for C-STS remains elusive as reported by Lin et al. (2024) who showed that decoder-only LLMs often underperform MLM-based embeddings in C-STS benchmarks. Interestingly, we find that encoding the condition given the sentence often outperforms the reverse setting in C-STS benchmarks. Moreover, we find that subtracting the embedding of the condition in a post-processing step further improves the performance across multiple LLM-based embedding models.
One disadvantage of using LLM-based embeddings compared to their MLM-based counterparts is the high dimensionality of the LLM-based embeddings. For example, embeddings produced by state-of-the-art (SoTA) LLM encoders such as NV-embed-v2 Lee et al. (2024) are 4096 dimensional, whereas MLM embeddings such as RoBERTa-base Liu et al. (2019) SimCSE Gao et al. (2021) embeddings are 768 dimensional. High dimensional embeddings are problematic both due to their high storage requirements and the high inner-product computation cost. Moreover, the embeddings obtained from LLMs are not necessarily aligned to the C-STS task because they are not fine-tuned on such tasks. To address these issue, we propose a supervised dimensionality reduction method that accurately projects LLM-based embeddings to low dimensional vector spaces, while also improving their performance in the C-STS task. We find that although popular unsupervised dimensionality reduction methods such as Principal Component Analysis (PCA) and Independent Component Analysis (ICA) are effective at reducing the dimensionality of the embedding space, they underperform our proposed supervised projection learning method. Further investigations into individual examples show that CASE increases the similarity between a sentence and the information emphasized by a given condition, while decreasing the same for information irrelevant to the condition as expected. We have submitted the source code and pre-processed data for reproducing our findings for anonymous reviewing, which will later be publicly released upon paper acceptance.
§ 2 Related Work
Conditional Semantic Textual Similarity:
Deshpande et al. (2023) proposed the C-STS task for measuring the similarity between two sentences under a given condition. They created a human-annotated dataset containing 18,908 instances where the semantic similarity between two sentences $s_{1}$ and $s_{2}$ is rated under two conditions $c_{\rm high}$ and $c_{\rm low}$ resulting in respectively high vs. low similarity between the two sentences. Moreover, they proposed cross-, bi- and tri-encoder baselines. Given a triplet $(s_{1},s_{2},c)$ a cross-encoder considers all interactions among tokens in $s_{1}$, $s_{2}$ and $c$. Although the cross-encoder setting benefits from having access to both $s_{1}$ and $s_{2}$ at the same time, it is computationally costly due to the large number of token interactions required for longer sentences and conditions. Moreover, it does not pre-compute conditional embeddings for the individual sentences, which is problematic when computing all pairwise similarities between sentences in a large set.
Bi-encoders overcome the limitations of cross-encoders by creating a condition-aware embedding for each sentence, and then compute C-STS using an efficient operation such as the inner-product. Tri-encoders separately encode $s_{1}$, $s_{2}$ and $c$, and then apply some late interactions between the condition’s and each sentence’s embeddings to compute C-STS. Despite the computational benefits gained by pre-encoding sentences and conditions separately, late interaction mechanisms for tri-encoders remain complex and under-perform in C-STS benchmarks.
Tu et al. (2024) found that the C-STS dataset created by Deshpande et al. (2023) contain ill-defined conditions and annotation errors, resulting in a significant discrepancy among the annotators for over half of the dataset. To address this, Tu et al. (2024) re-annotated the validation split from the original C-STS dataset. Moreover, they proposed a method to solve C-STS by first converting each condition into a question, and then using GPT-3.5 to extract the corresponding answers from the two sentences. Finally, C-STS is estimated as the cosine similarity between the SimCSE embeddings for the two answers. Although this QA formulation consistently improves the performance of all baseline encoders, it depends on multiple decoupled components such as converting conditions into questions, requiring a decoder LLM to extract the answer from each question, and using a separate encoder to compare the extracted answers, which increases the possibility of error propagation across components.
Yoo et al. (2024) proposed Hyper-CL, a contrastively learnt hypernetwork Ha et al. (2017) to selectively project the embeddings of $s_{1}$ and $s_{2}$ according to $c$. Hyper-CL follows a tri-encoder setting where $s_{1}$, $s_{2}$ and $c$ are first encoded separately using a sentence encoder. Next, a hypernetwork is trained to produce a linear transformation matrix conditioned on $c$. Finally, the embeddings of $s_{1}$ and $s_{2}$ are projected using this transformation matrix and their inner-product is computed. Hyper-CL improves the performance of tri-encoder models, but still under-performs bi-encoders for C-STS. Moreover, hypernetworks introduce an external parameter set that is three times larger than the SimCSE model used to encode each sentence, resulting in an excessively large memory space.
Lin et al. (2024) proposed a tri-encoder-based C-STS method where they used routers and heavy-light attention Ainslie et al. (2023) to select the relevant tokens to a given condition. Specifically, they used the query vector of the [CLS] token of the condition to compute attention scores for the tokens in the sentence, which are subsequently used to compute a sentence embedding. Their method outperforms Hyper-CL for the tri-encoder setting.
Li et al. (2024) proposed a cross-encoder approach which predicts C-STS scores, without creating conditional embeddings. Prior work on C-STS has shown cross-encoders to perform poorly despite having access to both sentences simultaneously. Li et al. (2024) showed that this is due to the irrelevant information in the two sentences to the condition and proposed a token re-weighting strategy, inspired by object detection in computer vision Shi et al. (2023); Jaegle et al. (2021). Concretely, they compute two cross-attention matrices between $(s_{1},s_{2})$ and $c$, which are subsequently used to compute the correlations for the sentence or condition tokens. Although their method improves the performance of cross-encoder-based C-STS measurement, it still underperforms bi-encoders. Consequently, we propose bi-encoders for C-STS, which are computationally efficient and accurate.
Figure 2: Overview of CASE. An LLM is prompted with $I(s)$ to obtain two separate embeddings $f(c;I(s_{1}))$ and $f(c;I(s_{2}))$ for the condition $c$ for two sentences $s_{1}$ and $s_{2}$. The unconditional embedding $f(c;I(\emptyset))$ is then computed using the prompt $I(\emptyset)$ and subtracted from each of those embeddings. Finally, the embeddings are projected to a lower-dimensional space using a learnt projection $g$ and their cosine similarity is computed.
Text Embeddings from LLMs:
MLMs use a bi-directional attention Devlin et al. (2018), where the information from tokens appearing in positions both before and after the current token are used to predict the embedding for the current target token in the sequence that must be encoded. In contrast, decoder-only LLMs are trained with a causal attention mask Vaswani et al. (2017), which prevents information leakage from future tokens by allowing the decoder to attend only to previous positions during auto-regressive text generation. Although this makes sense for decoders, it is sub-optimal when using LLMs for encoding a given sequence as evident from the poor performance of GPT models Lin et al. (2024) compared to similar-sized BERT Devlin et al. (2018) or T5 Chung et al. (2024) models on various natural language understanding benchmarks Wang et al. (2019).
BehnamGhader et al. (2024) proposed a post-processing method for obtaining text embeddings from decoder-only LLMs following three steps: (a) enable bidirectional attention to overcome the restrictions due to causal attention, (b) train the model to predict the masked next token using bi-directional attention, and (c) use unsupervised contrastive learning to compute better sequence representations. Finally, a pooling method is applied on the token embedding sequence to create a fixed-dimensional embedding for the input text such as the embedding of the last token in the sequence (i.e. last token pooling) Meng et al. (2024); Li et al. (2023) or the average over all token embeddings (i.e. mean pooling) Wang et al. (2024). Moreover, Lee et al. (2024) proposed a latent attention layer where they compute the cross attention between the last hidden layer from the decoder and a trainable latent array to compute a weighted pooling mechanism.
§ 3 Condition-Aware Sentence Embeddings
An overview of our proposed method is shown in Figure 2, which consists of two-steps. In the first step (§ 3.1), we create two separate embeddings for the condition considering each of the two sentences, one at a time, in the instruction prompt shown to an LLM. Note that it is the condition that is being encoded and the tokens in the sentence (similar to all other tokens in the instruction) are simply modifying the attention scores computed for the tokens in the condition. Intuitively, it can be seen as each sentence filling some missing information asked in the condition.
Embeddings obtained from LLMs are typically high dimensional and not necessarily fine-tuned for the C-STS task. Therefore, in the second step (§ 3.2), we learn a supervised projection using the training split from the C-STS dataset to project each of the two condition embeddings to a lower-dimensional vector space. Finally, the C-STS between two sentences is computed as the cosine similarity between the corresponding projected embeddings.
§ 3.1 Extracting Embeddings from LLMs
Given an LLM-based encoder, $f$, we create a $d$-dimensional embedding $f(c;I(s))\in\mathbb{R}^{d}$ for a condition $c$, given the sentence $s$. Here, $I$ is an instruction template that takes $c$ as an argument. We use the following prompt template as $I(s)$Retrieve semantically similar texts to the Condition, given the Sentence : [SENTENCE], where we substitute $s$ in the placeholder [SENTENCE]. Next, we provide $c$ as the input text to be encoded by the LLM following the instruction $I(s)$. Finally, the token embeddings of $c$ are aggregated according to one of the pooling methods described previously to create $f(c;I(s))$. Although our focus here is to create condition-aware embeddings, it is noteworthy that we can also obtain unconditional embeddings for a sentence by dropping the condition in the above prompt. Specifically, we use the prompt $I(\emptyset)$ Retrieve semantically similar texts to a given Sentence for this purpose, and denote this unconditional embedding of $c$ by $f(c;I(\emptyset))$. As we see later in our experiments, by subtracting $f(c;I(\emptyset))$ from $f(c;I(s))$, we can reduce the effect of tokens in the condition that are irrelevant to the sentence, thereby consistently improving accuracy of the condition-aware embeddings. This first step is fully unsupervised and a zero-shot prompt template is used as $I$.
Recall that both $s$ and $c$ are text strings, and it is possible to swap the sentence and condition in the above formulation to obtain an embedding $f(s;I(c))$ for the sentence, given the condition. However, as shown later in our experiments, comparing the embedding for $c$ created under $s_{1}$ and $s_{2}$ results in better performance on the C-STS benchmark for all LLM encoders. This is because $s_{1}$ and $s_{2}$ contain many irrelevant information to $c$, which will affect the cosine similarity computed between $f(s_{1};I(c))$ and $f(s_{2};I(c))$. On the other hand, the cosine similarity between $f(c;I(s_{1}))$ and $f(c;I(s_{2}))$ is a more accurate estimate of the C-STS between $s_{1}$ and $s_{2}$ under $c$ because it is purely based on the meaning alterations to $c$ under $s_{1}$ and $s_{2}$ separately.
§ 3.2 Supervised Projection Learning
The LLM-based embeddings computed in § 3.1 has two main drawbacks. First, relative to MLMs-based sentence embeddings, LLMs produce much higher dimensional embeddings, which can be problematic due to their memory requirements (especially when operating on a limited GPU memory) and the computational cost involved in inner-product computations. In tasks such as dense retrieval, we must compare millions of documents against a query to find the nearest neighbours under strict latency requirements, and low-dimensional embeddings are preferable. Second, although LLMs are typically trained on massive text collections and instruction-tuned for diverse tasks Muennighoff et al. (2022), their performance on C-STS tasks have been poor Lin et al. (2024). As seen from our condition-aware prompt template, an LLM must be able to separately handle a variable condition statement and a fixed instruction. This setup is different from most tasks on which LLMs are typically trained, where the instruction remains fixed across all inputs. Therefore, it is important to perform a supervised fine-tuning step to LLM embeddings before they are used for C-STS tasks.
To address the above-mentioned drawbacks, we propose a supervised projection learning method. Specifically, we freeze the model parameters of the LLM and use an Multi-layer Perception (MLP) layer that takes $f(c;I(s))$ as the input and returns a $k$-dimensional ($k\leq d)$ embedding $g(f(c;I(s);\theta)$, where $\theta$ denotes the learnable parameters of the MLP. Finally, we define ${\rm CASE}(s,c)$, as the projection of the offset between the conditional and the unconditional embeddings of $c$ under $s$, given by (1).
${\rm CASE}(s,c)=g(f(c;I(s_{1}))-f(c;I(\emptyset));\theta)$ (1)
We use the human similarity ratings $r$ in the C-STS train instances $\mathcal{D}$ to learn $\theta$. Specifically, we minimise the squared error between the human ratings and the cosine similarity computed using the corresponding CASE as given by (2).
$\sum_{(s_{1},s_{2},c,r)\in\mathcal{D}}\left(\cos({\rm CASE}(s_{1},c),{\rm CASE}(s_{2},c))-r\right)^{2}$ (2)
Here, $\cos$ denotes the cosine similarity between the projected embeddings. We use Adam optimiser Kingma and Ba (2014) to find the optimal $\theta$ that minimises the loss given by (2). Recall that only the MLP parameters are updated during this projection learning step, while keeping the parameters of the LLM fixed, which makes it extremely efficient. For example, it takes less than 5 minutes to learn this projection even for the largest (4096 dimensional) embedding spaces using the train split of the C-STS dataset. Using the learnt projection, we compute the C-STS between $s_{1}$ and $s_{2}$ under $c$ as the cosine similarity between ${\rm CASE}(s_{1},c)$ and ${\rm CASE}(s_{2},c)$.
§ 4 Experiments and Results
To evaluate the effectiveness of the LLM-based and MLM-based sentence embeddings as described in § 3, we apply six sentence encoders, out of which three are LLM-based: NV-Embed-v2 (4096 dimensional NV), SFR-Embedding-Mistral (4096 dimensional SFR), gte-Qwen2-7B-instruct (3584 dimensional GTE) and three are MLM-based: Multilingual-E5-large-instruct (1024 dimensional E5), sup-simcse-roberta-large (1024 dimensional SimCSE_large), and sup-simcse-bert-base-uncased (768 dimensional SimCSE_base). Further details of these models are provided in Appendix A.
We evaluate model performance on different pooling methods, prompt settings, and sentence constructions. Then, we evaluate the dimensionality reduction methods under supervised and unsupervised settings to learn the projection for CASE (PCA, ICA, and linear and non-linear MLPs). Both the linear and non-linear MLPs are Siamese bi-encoders with weight-sharing. As explained in § 3.2, they take $f(c;I(s_{1}))$ and $f(c;I(s_{2}))$ as the input embeddings, and return the projected embeddings ${\rm CASE}(s_{1},c)$ and ${\rm CASE}(s_{2},c)$ as the outputs, which can be compared using a similarity metric such as the cosine similarity.
Our linear MLP is a single-layer linear transformation:
$\mathbf{z}=\operatorname{Dropout}\left(\text{\bf W}\bm{e}\right)$ (3)
where $\text{\bf W}\in\mathbb{R}^{d^{\prime}\times d}$ is the learned projection matrix. A dropout layer is applied to reduce any overfitting Hinton et al. (2012).
Our non-linear MLP is a two-layer MLP:
$\mathbf{h}$ $=\operatorname{Dropout}\left(\operatorname{ReLU}\left(\text{\bf W}_{1}\bm{e}\right)\right)$ (4)
$\mathbf{z}$ $=\operatorname{Dropout}\left(\operatorname{ReLU}\left(\text{\bf W}_{2}\bm{h}\right)\right)$ (5)
Hyperparameters are tuned on a held-out validation set. We select a learning rate of $10^{-3}$ and a batch size of 512. The dropout rate is set to $20\%$ for linear MLP and $15\%$ for non-linear MLP.
To address annotation errors in the original C-STS dataset such as ambiguous and invalid conditions, Tu et al. (2024) re-annotated the original validation set. To conduct a more accurate and reliable evaluation, we use the original C-STS train set and re-annotated validation set. A 70-30 split is used for the re-annotated validation set, with randomly selected 70% of the data allocated for validation and the remaining 30% for testing. We use a single p3.24xl EC2 instance (8x V100 GPUs) for learning sentence embeddings, and a separate NVIDIA RTX A6000 GPU for supervised projection learning. Scikit-learn 1.3.0 is used for PCA and ICA. Torch 2.0.1+cu117 is used for MLP proejction. These settings are fixed across all experiments. In the case of reducing 4096-dimensional LLM-based sentence embeddings to 512-dimensional, training a linear MLP for CASE takes approximately 1.5 minutes, whereas a non-linear MLP requires about 5 minutes (wall-clock time).
§ 4.1 Evaluation Metrics
We evaluate the performance of sentence embedding models on two metrics: Spearman Rank correlation and Accuracy.
Spearman Rank Correlation:
We compute Spearman’s rank Correlation between the similarity scores by CASE and the re-annotated human ratings on the test set.
Accuracy
Spearman correlation is highly sensitive to small variations in similarity scores, which can be affected by the noise in human annotations. Given the subjectivity and ambiguity of human ratings, we introduce Accuracy as a more robust alternative to assess whether the model correctly captures the relative impact of conditions on semantic similarity. Different from Spearman correlation, which considers actual similarity values, accuracy only evaluates whether the predicted similarity under the higher-rated condition $c_{\rm high}$ is greater than that under the lower-rated condition $c_{\rm low}$.
For each sentence pair $\left(s_{1},s_{2}\right)$, there exist two conditions $c_{1}$ and $c_{2}$ with corresponding human labelled similarity scores $y_{1}$ and $y_{2}$, where $y_{1}>y_{2}$.
Cosine similarity is computed between CASE under the same condition for the similarity score
$\operatorname{sim}_{c_{1}}=\operatorname{cos}\left({\rm CASE}(s_{1},c_{1}),{\rm CASE}(s_{2},c_{1})\right)$ (6)
$\operatorname{sim}_{c_{2}}=\operatorname{cos}\left({\rm CASE}(s_{1},c_{2}),{\rm CASE}(s_{2},c_{2})\right)$ (7)
A prediction is considered correct if
$\left(\operatorname{sim}_{c_{1}}-\operatorname{sim}_{c_{2}}\right)\left(y_{1}-y_{2}\right)>0,$ (8)
which evaluates whether the model’s predicted similarity ranking aligns with the human annotations. Then, the accuracy is given by (9)
$\frac{\sum_{i=1}^{N}1\left[\left(\operatorname{sim}_{c_{1}}^{(i)}-\operatorname{sim}_{c_{2}}^{(i)}\right)\left(y_{1}^{(i)}-y_{2}^{(i)}\right)>0\right]}{N},$ (9)
where $N$ is the total number of test instances, and $1[\cdot]$ is the indicator function.
§ 4.2 C-STS Measurement
To generate CASE, we apply different ways to construct the prompt for LLM-based embeddings and to concatenate the condition and sentence for MLM-based embeddings. For LLM-based models, we have two main settings: (a) $\operatorname{cond}=f(c;I(s))$, where we encode the condition given the sentence, and (b) $\operatorname{sent}=f(s;I(c))$, where we encode the sentence given the condition as explained in § 3.1. For MLM-based models, we evaluate the two settings: (a) $\operatorname{CONC}(c+s)$, where we concatenate sentence after the condition, and (b) $\operatorname{CONC}(s+c)$, where we concatenate condition after the sentence. For each setting, we further evaluate the effect of subtracting the embedding for the condition, $\bm{c}=f(c;I(\emptyset))$.
The test performance for different settings and models are shown in Table 1. For LLM-based embeddings, cond consistently reports higher Spearman correlation than sent, suggesting that embedding the condition given the sentence is more effective for C-STS measurement than embedding the sentence given the condition. Moreover, we see that subtracting $\bm{c}$ further improves performance both in terms of Spearman correlation and accuracy across all settings, except for accuracy in SFR. The former approach reduces the noise due to the tokens in a sentence, which are irrelevant to the given condition. For MLM-based embeddings, subtracting $\bm{c}$ also improves performance. It reduces condition-specific information that are not altered by the sentence, thus allowing CASE to focus on the information that varies between the two sentences being compared. Moreover, we discovered that subtracting $\bm{c}$ in a post-processing step improves isotropy of the embeddings as shown in Appendix B. This is in-line with prior work reporting a correlation between isotropy and improved performance in embedding models Rajaee and Pilehvar (2022); Su et al. (2021). The effect of pooling method on performance is discussed in Appendix C, where we find the latent pooling in NV to perform best. Therefore, we use the ($\operatorname{cond}-\bm{c}$) setting (which reports the best performance) for the six sentence encoders to conduct the subsequent experiments.
Model sent/cond? Spear. Acc.
NV sent - c 16.98 52.24
sent 22.07 59.89
cond - c 31.32 64.91
cond 27.02 48.02
SFR sent - c 19.54 57.52
sent 11.89 43.01
cond - c 20.38 52.51
cond 18.32 46.44
GTE sent - c 7.16 42.48
sent 7.16 37.20
cond - c 20.40 54.08
cond 16.58 45.12
E5 sent - c 11.08 42.21
sent 3.77 33.77
cond - c 12.01 42.74
cond 6.18 37.47
SimCSE_large $\operatorname{CONC}(s+c)$ 5.59 37.46
$\operatorname{CONC}(c+s)$ 4.00 35.09
$\operatorname{CONC}(s+c)$ c 8.32 34.56
$\operatorname{\textbf{CONC}}\bm{(c+s)}$ c 8.58 48.02
SimCSE_base $\operatorname{CONC}(s+c)$ 4.37 39.81
$\operatorname{CONC}(c+s)$ 1.25 34.82
$\operatorname{\textbf{CONC}}\bm{(s+c)}$ c 7.05 42.74
$\operatorname{CONC}(c+s)$ c 6.00 36.67
Table 1: Spearman and accuracy scores for different sentence embedding models and encoding settings.
(a) Test Spearman for linear MLP
(b) Test Spearman for non-linear MLP
Figure 3: Spearman correlation on test set for different models over training steps with dimensionality 512. The $y$-axes of both subfigures are aligned, facilitating a direct comparison of the Spearman correlation coefficients across the two line charts, with the same colour for the same model. Best viewed in colour.
Figure 4: Spearman correlation coefficients of the four dimensionality reduction methods on the test set for NV-Embed-v2 (NV) over different dimensionalities.
Model Non-linear MLP Linear MLP PCA ICA
NV 69.08 69.95 32.04 19.13
SFR 62.85 59.22 19.86 -1.99
GTE 64.16 56.10 22.82 7.26
E5 62.12 47.03 11.43 -0.17
SimCSE_large 56.67 45.96 8.22 -4.89
SimCSE_base 56.60 39.54 6.19 -1.66
Table 2: Spearman correlation of embedding models based on supervised and unsupervised dimensionality reduction methods with reduced dimensionality $512$.
s1: Young woman in orange dress about to serve in tennis game, on blue court with green sides.
s2: A girl playing tennis wears a gray uniform and holds her black racket behind her.
cos(s1, s2)   $0.5006\rightarrow 0.9016$
Condition 1: The color of the dress.  Rating: 1 Condition 2: The name of the game.  Rating: 5
cos(s1, s2; c1)   $0.4757\rightarrow 0.2522$ cos(s1, s2; c2)   $0.6233\rightarrow 0.9660$
s1: Two snow skiers with ski poles and snow skis, standing on top of a snow covered mountain with other skiers around them.
s2: A skier stands alone at the top of a snowy slope with blue skies and mountains in the distance.
cos(s1, s2)   $0.6318\rightarrow 0.7702$
Condition 1: The number of person.  Rating: 1 Condition 2: The type of job.        Rating: 5
cos(s1, s2; c1)    $0.6358\rightarrow 0.2788$ cos(s1, s2; c2)$0.7502\rightarrow 0.9539$
s1: A bunch of people standing around at the beach with a kite in the air.
s2: a beach scene with a beach chair decorated with the Canadian Flag and surfers walking by with their surfboards
cos(s1, s2)   $0.3988\rightarrow 0.5350$
Condition 1: The type of hobby.   Rating: 1 Condition 2: The type of location.      Rating: 5
cos(s1, s2; c1)   $0.4386\rightarrow 0.4886$ cos(s1, s2; c2)$0.4825\rightarrow 0.8582$
Table 3: Cosine similarity scores ($\cos$) for two sentences under different conditions are computed using NV embeddings (shown on the left side of the arrows) vs. using MLP projected $512$ embeddings (i.e. CASE embeddings given by (1)) (shown on the right side of the arrows). Compared to the unconditional similarity between the two sentences, conditional similarity scores computed using NV embeddings align well with the human ratings. Moreover, the conditional similarities are further appropriately amplified by the supervised projection.
We show the training curves for our supervised MLPs in Figure 3. Overall, the non-linear MLPs achieve significantly higher Spearman correlation than the linear MLPs, except for NV, where the linear MLP performs slightly better than the non-linear MLP. Moreover, non-linear MLPs converge faster, typically reaching their peak performance within $20$ epochs, after which the performance declines due to overfitting. The performance of the linear MLPs, gradually increases and eventually converges as the training progresses. NV consistently performs the best for both linear and non-linear MLPs.
To explore the relationship between test performance and dimensionality, we use NV to evaluate our supervised dimensionality reduction methods (MLPs) as well as the unsupervised methods (PCA and ICA), as shown in Figure 4. Overall, the supervised MLPs achieve much higher Spearman correlations while significantly reducing the dimensionality. Their performance stabilises at relatively high levels as the dimensionality exceeds $200$ and peaks around $500$. In this way, we achieve an 8×8$\times$8 × compression of the sentence embeddings (original dimensionality $4096$) while maintaining high performance. In contrast, unsupervised PCA and ICA do not effectively capture the conditional semantics in sentence embeddings. While PCA preserves the performance of the original sentence embeddings, it does not improve the performance on the C-STS task. ICA, on the other hand, performs even worse, showing a consistent decline in performance as dimensionality increases, which could be attributable to the lack of sufficiently independent components in sentence embeddings.
We compare the different sentence encoders in Table 2 for a fixed (i.e. 512) dimensional projection from their original embeddings using Spearman correlation. NV obtains the highest Spearman correlation across all sentence encoders and dimensionality reduction methods. GTE and SFR also achieve high Spearman correlation coefficients under the non-linear MLPs. SimCSE_large and SimCSE_base have the lowest Spearman correlation coefficients, regardless of the dimensionality reduction method being used. Overall, we see that LLM-based embeddings perform better than the MLM-based embeddings on the C-STS task.
Similarity scores computed for three sentence pairs are shown in Table 3 for the original high-dimensional LLM-based embeddings and low-dimensional projected CASE. By comparing the similarity scores with and without the supervised projection, we can evaluate the ability of CASE to focus on the condition-related information. Compared to the unconditional similarity between the two sentences, conditional similarities scores computed using NV embeddings align well with the human ratings. For example, in the top row in Table 3, we see that the unconditional similarity between the two sentences reduces from $0.5006$ to $0.4747$ under $c_{1}$, while increasing to $0.6233$ under $c_{2}$. Moreover, the conditional similarities are further appropriately amplified by CASE using the supervised projection (i.e. decreasing to $0.2522$ under $c_{1}$, while increasing to $0.9660$ under $c_{2}$). Specifically, CASE reduces the similarity between the two sentences under the lower-rated condition, while increasing the same under the high-rated condition. This demonstrates that CASE is able to effectively capture the shift in conditional meaning between sentences under different conditions.
§ 5 Conclusion
We propose CASE in a novel setting, which encodes the condition under sentence by subtracting the condition as a post-processing step when measuring C-STS. Our findings show that LLM-based sentence embeddings consistently outperform MLM-based embeddings for C-STS. Moreover, we introduce an efficient supervised projection learning method to reduce the dimensionality of LLM-based embeddings, while improving the performance in C-STS. CASE correctly compares the tokens relevant to a given condition between the two sentences as shown in further qualitative results in Appendix D.
§ 6 Limitations
We use the original C-STS training set and the re-annotated validation set in this paper. The dataset, especially the training set, includes annotation errors, such as invalid conditions and subjective human annotations. These disadvantages hurt the model’s ability to capture the semantics and perform the C-STS task. This further introduces ambiguity and inaccuracy to CASE. Manually re-annotating the training set, which contains 11342 instances, is costly and is deferred to future work.
Our evaluations cover only English, which is a morphologically limited language. To the best of our knowledge, C-STS datasets have not been annotated for languages other than English, which has limited all prior work on C-STS to conduct experiments using only English data. However, the sentence encoders we used in our experiments support multiple languages. Therefore, we consider it to be an important future research direction to create multilingual datasets for C-STS and evaluate the effectiveness of our proposed method in multilingual settings.
There is a large number of sentence encodes (over 240 models as at March 2025 evaluated on Massive Text Embedding Benchmark (MTEBMuennighoff et al. (2023) leaderboard [^1] https://huggingface.co/spaces/mteb/leaderboard). However, due to computational limitations, we had to select a subset covering the best performing (top ranked on MTEB at the time of writing) models for our evaluations. We will release our source code and the evaluation framework such that the research community can evaluate our proposed method with any sentence encoder for C-STS.
§ 7 Ethical Concerns
We did not collect or annotate any datasets in this project. Instead, we use existing C-STS datasets annotated and made available by Deshpande et al. (2023) and Tu et al. (2024). To the best of our knowledge, no ethical issues have been raised regarding those datasets.
We use multiple pre-trained and publicly available MLM- and LLM-based sentence encoders Kaneko and Bollegala (2021). Both MLMs and LLMs are known to encode unfair social biases such as gender or racial biases. We have not evaluated how such social biases would be influenced by the CASE learning method proposed in this work. Therefore, we consider it would be important to measure the social biases in CASE created in this work before they are deployed in real-world applications.
Supplementary Materials
Appendix A Models
To evaluate the effectiveness of the LLM-based and MLM-based sentence embeddings as described in § 3, we apply six sentence encoders, out of which three are LLM-based: NV-Embed-v2 (4096 dimensional and uses latent pooling) (NV) [^2] https://huggingface.co/nvidia/NV-Embed-v2, SFR-Embedding-Mistral (4096 dimensional and uses average pooling) (SFR) [^3] https://huggingface.co/Salesforce/SFR-Embedding-Mistral, gte-Qwen2-7B-instruct (3584 dimensional and uses last token pooling) (GTE) [^4] https://huggingface.co/Alibaba-NLP/gte-Qwen2-7B-instruct and three are MLM-based: Multilingual-E5-large-instruct (1024 dimensional and uses average pooling) (E5) [^5] https://huggingface.co/intfloat/multilingual-e5-large-instruct, sup-simcse-roberta-large (1024 dimensional and uses last token pooling) (SimCSE_large) [^6] https://huggingface.co/princeton-nlp/sup-simcse-roberta-large, and sup-simcse-bert-base-uncased (786 dimensional and uses last token pooling) (SimCSE_base) [^7] https://huggingface.co/princeton-nlp/sup-simcse-bert-base-uncased.
Appendix B Isotropy for Embeddings
We first use the embedding-to-mean cosine similarity distribution to measure the isotropy of the embeddings. Given a set of embeddings $S=\{\bm{x}_{1},\bm{x}_{2},\dots,\bm{x}_{n}\}$, we first compute the mean embedding vector $\bm{\mu}=\frac{1}{n}\sum_{i=1}^{n}\bm{x}_{i}$. Then, for each embedding $\bm{x}_{i}\in S$, we compute its cosine similarity with the mean vector $\mu$, i.e., $\cos(\bm{x}_{i},\bm{\mu})=\frac{\bm{x}_{i}{}^{\top}\bm{\mu}}{\left|\left|\bm{x}_{i}\right|\right|\left|\left|\bm{\mu}\right|\right|}$. The distribution of these embedding-to-mean cosine similarities is then analysed to characterise the embedding space – a distribution sharply peaked near 1 indicates anisotropy, whereas a broader, more uniform distribution suggests a more isotropic geometry.
From Table 4, Figure 5, and Figure 6, we see that the embeddings after subtracting c have a lower mean cosine similarity to the mean vector and a higher standard deviation, indicating that they are more spread out in the embedding space. In contrast, the embeddings without subtracting c are more clustered around a central direction (higher mean, lower standard deviation), reflecting anisotropy, a tendency for vectors to concentrate in a narrow region. Therefore, embeddings after subtracting c tend to be more isotropic, indicating better distributional diversity.
(a) NV: cond - c
(b) NV: cond
(c) SFR: cond - c
(d) SFR: cond
(e) GTE: cond - c
(f) GTE: cond
Figure 5: Embeddings-to-mean cosine similarity distributions across three LLM-based models. Each row compares cond - c and cond representations.
(a) E5: cond - c
(b) E5: cond
(c) SimCSE_base: $\operatorname{CONC}(c+s)$ c
(d) SimCSE_base: $\operatorname{CONC}(c+s)$
(e) SimCSE_large: $\operatorname{CONC}(c+s)$ c
(f) SimCSE_large: $\operatorname{CONC}(c+s)$
Figure 6: Embeddings-to-mean cosine similarity distributions across three MLM-based models. Each row compares two embedding types.
Model Embedding Type Mean Std
NV cond - c 0.407 0.084
cond 0.492 0.067
SFR cond - c 0.537 0.055
cond 0.708 0.021
GTE cond - c 0.489 0.067
cond 0.696 0.036
E5 cond - c 0.542 0.056
cond 0.897 0.010
SimCSE_base $\operatorname{CONC}(c+s)$ c 0.254 0.095
$\operatorname{CONC}(c+s)$ 0.347 0.088
SimCSE_large $\operatorname{CONC}(c+s)$ c 0.248 0.110
$\operatorname{CONC}(c+s)$ 0.333 0.085
Table 4: Cosine similarity to mean vector: comparing mean and standard deviation of two embedding types across three LLM-based and three MLM-based models.
Another method to measure the isotropy is Principle Components (IPC) Mu et al. (2018). To address the numerical instability issues when using PCA for isotropy measurement in high-dimensional embedding spaces, we use an approximation method for computing the IPC. Given a normalized embedding matrix $E\in\mathbb{R}^{n\times d}$ where each row represents a unit vector, instead of computing the eigenvectors of the covariance matrix, we randomly sample $k$ unit vectors ${u_{1},u_{2},...,u_{k}}$ from the unit hypersphere $\mathbb{S}^{d-1}$. For each sampled direction $u_{i}$, we compute the function $F(u_{i})=\sum_{j=1}^{n}\exp(E_{j}\cdot u_{i})$, where $E_{j}$ represents the $j$-th embedding vector. The approximated IPC is then defined as the ratio between the minimum and maximum values of $F$: $\text{IPC}{\text{approx}}=\frac{\min{i\in{1,...,k}}F(u_{i})}{\max_{i\in{1,...,k}}F(u_{i})}$. If the approximated IPC value is close to 1, it means the embedding space shows high isotropy where vectors are uniformly distributed across all directions in the high-dimensional space instead of clustering in some dominant directions. Conversely, when the IPC value approaches 0, it means significant anisotropy in the embedding space. We use $k=1000$ random directions to compute the approximation of IPC. Table 5 shows that the post-processing step of subtracting the condition c gives higher approximated IPC values, indicating the improvement of isotropy.
Model IPC (-c) IPC
NV 0.9611 0.9471
SFR 0.9463 0.9266
GTE 0.9490 0.9310
E5 0.8906 0.8368
SimCSE_base 0.9461 0.9297
SimCSE_large 0.9499 0.9406
Table 5: Approximated IPC values for each model. The left column of IPC (-c) lists values with post-processing step of subtracting the condition c.
Appendix C Full Results for all Models
Each LLM encoder has its own preferred pooling method, recommended by the authors of those models. From Table 1 we see that the performance varies significantly depending on the pooling method being used, while the latent attention pooling used in NV reporting the best results. Note that NV does not support last or average pooling, while latent pooling is not supported by the other models.
Model Pooling sent/cond? Spear. Acc.
NV latent sent - c 16.98 52.24
latent sent 22.07 59.89
latent cond - c 31.32 64.91
latent cond 27.02 48.02
SFR last sent - c 11.88 46.97
last sent -0.18 33.77
last cond - c 19.28 50.66
last cond 13.00 44.33
average sent - c 19.54 57.52
average sent 11.89 43.01
average cond - c 20.38 52.51
average cond 18.32 46.44
average sent - c 19.54 57.52
average sent 11.89 43.01
average cond - c 20.38 52.51
average cond 18.32 46.44
GTE last sent - c 7.16 42.48
last sent 7.16 37.20
last cond - c 20.40 54.08
last cond 16.58 45.12
average sent - c 13.01 42.48
average sent 11.34 36.67
average cond - c 17.87 42.22
average cond 18.42 42.22
E5 average sent - c 11.08 42.21
average sent 3.77 33.77
average cond - c 12.01 42.74
average cond 6.18 37.47
last sent - c 9.28 41.69
last sent 0.49 31.13
last cond - c 8.90 39.84
last cond 3.05 34.56
SimCSE_large $\operatorname{CONC}(s+c)$ 5.59 37.46
$\operatorname{CONC}(c+s)$ 4.00 35.09
$\operatorname{CONC}(s+c)$ c 8.32 34.56
$\operatorname{\textbf{CONC}}\bm{(c+s)}$ c 8.58 48.02
SimCSE_base $\operatorname{CONC}(s+c)$ 4.37 39.81
$\operatorname{CONC}(c+s)$ 1.25 34.82
$\operatorname{\textbf{CONC}}\bm{(s+c)}$ c 7.05 42.74
$\operatorname{CONC}(c+s)$ c 6.00 36.67
Table 6: Spearman and accuracy scores for sentence embedding models. The original dimensionality of each model is indicated in parentheses. latent denotes latent attention pooling for NV, whereas last and average correspond to last token pooling and average pooling, respectively.
Appendix D Additional Qualitative Analysis
s1: Young woman in orange dress about to serve in tennis game, on blue court with green sides.
s2: A girl playing tennis wears a gray uniform and holds her black racket behind her.
cos(s1, s2)   $0.5006\rightarrow 0.9016$
Condition 1: The color of the dress. Condition 2: The name of the game.
Answer 1: orangeAnswer 2: grayRating: 1 Answer 1: tennisAnswer 2: tennisRating: 5
cos(s1, s2; c1)   $0.4757\rightarrow 0.2522$ cos(s1, s2; c2)   $0.6233\rightarrow 0.9660$
cos(s1, orange; c1)   $0.1986\rightarrow 0.3551$ cos(s1, tennis; c2)   $0.1135\rightarrow 0.6448$
cos(s2, gray; c1)   $0.0559\rightarrow 0.6061$ cos(s2, tennis; c2)   $0.0983\rightarrow 0.6426$
s1: Two snow skiers with ski poles and snow skis, standing on top of a snow covered mountain with other skiers around them.
s2: A skier stands alone at the top of a snowy slope with blue skies and mountains in the distance.
cos(s1, s2)   $0.6318\rightarrow 0.7702$
Condition 1: The number of person. Condition 2: The type of job.
Answer 1: twoAnswer 2: oneRating: 1 Answer 1: skierAnswer 2: skierRating: 5
cos(s1, s2; c1)   $0.6358\rightarrow 0.2788$ cos(s1, s2; c2)   $0.7502\rightarrow 0.9539$
cos(s1, two; c1)   $-0.0970\rightarrow 0.7014$ cos(s1, skier; c2)   $0.2488\rightarrow 0.8016$
cos(s2, one; c1)   $0.0227\rightarrow 0.7953$ cos(s2, skier; c2)   $0.3409\rightarrow 0.8032$
s1: A bunch of people standing around at the beach with a kite in the air.
s2: a beach scene with a beach chair decorated with the Canadian Flag and surfers walking by with their surfboards
cos(s1, s2)   $0.3988\rightarrow 0.5350$
Condition 1: The type of hobby. Condition 2: The type of location.
Answer 1: kite flyingAnswer 2: surfRating: 1 Answer 1: beachAnswer 2: beachRating: 5
cos(s1, s2; c1)   $0.4386\rightarrow 0.4886$ cos(s1, s2; c2)   $0.4825\rightarrow 0.8582$
cos(s1, kite flying; c1)   $0.3457\rightarrow 0.7717$ cos(s1, beach; c2)   $0.2254\rightarrow 0.7281$
cos(s2, surf; c1)   $0.1034\rightarrow 0.6665$ cos(s2, beach; c2)   $0.1129\rightarrow 0.6703$
Table 7: Example of similarity scores for two conditions applied to the same sentence pair, based on linear MLP with dimensionality $512$ on NV. The table shows how supervised MLP projection improves the CASE for C-STS task. cos(,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) denotes cosine similarity. Answer 1 and Answer 2 refer to the corresponding answers for Sentence 1 and Sentence 2 under conditions, respectively. The predicted similarity scores before and after applying supervised MLP are listed on the left and right of the arrow. That is, the similarity score for original high-dimensional LLM-based embeddings is on the left of the arrow, while the similarity score for CASE is on the right.
We briefly show an example of our pipeline in Table 7. We treat the conditions as questions for sentences and extract the relevant information as answers to compare whether CASE can focus on the condition-related information. The overall similarity trend is in consistent with the actual ratings. For all sentence and answer pairs, the similarity scores increase after supervised projection. This demonstrates that the embeddings effectively capture the conditional meaning shift with higher condition-dependent similarity.