### v4

# CASE – Condition-Aware Sentence Embeddings for Conditional Semantic Textual Similarity Measurement

Abstract
The meaning conveyed by a sentence often depends on the context in which it appears. Despite the progress of sentence embedding methods, it remains unclear as how to best modify a sentence embedding conditioned on its context. To address this problem, we propose Condition-Aware Sentence Embeddings (CASE), an efficient and accurate method to create an embedding for a sentence under a given condition. First, CASE creates an embedding for the condition using a Large Language Model (LLM) encoder, where the sentence influences the attention scores computed for the tokens in the condition during pooling. Next, a supervised method is learnt to align the LLM-based text embeddings with the Conditional Semantic Textual Similarity (C-STS) task. We find that subtracting the condition embedding consistently improves the C-STS performance of LLM-based text embeddings by improving the isotropy of the embedding space. Moreover, our supervised projection method significantly improves the performance of LLM-based embeddings despite requiring a small number of embedding dimensions.111code: https://github.com/LivNLP/CASE
CASE – Condition-Aware Sentence Embeddings for
Conditional Semantic Textual Similarity Measurement
Gaifan Zhang1   Yi Zhou2   Danushka Bollegala1,3 1 University of Liverpool  2 Cardiff University  3 Amazon {sggzhan8,danushka}@liverpool.ac.uk, zhouy131@cardiff.ac.uk
§ 1 Introduction
Measuring Semantic Textual Similarity (STS) between sentences is a fundamental task in NLP Majumder et al. (2016). It is also important in training the sentence encoders Reimers and Gurevych (2019). Recent approaches, such as contrastive learning, use semantic similarity as an objective to improve the quality of sentence embeddings Gao et al. (2021). However, measuring sentence similarity is a complex task, which depends on the aspects being considered in the sentences being compared. To address this problem, Deshpande et al. (2023) proposed the C-STS task along with a human-annotated dataset. They designed the C-STS task and the dataset by assigning different conditions (semantic aspects) to each pair of sentences ($s_{1}$, $s_{2}$), a condition of high similarity $c_{\rm high}$ and a condition of low similarity $c_{\rm low}$, resulting in different similarity ratings for the same pair of sentences. This reduces subjectivity and ambiguity in the measurement of similarity between two sentences. As shown in Figure 1, human annotators are required to assign different similarity scores under different conditions, rather than a single score as in the traditional STS task. Many real-world applications can be seen as C-STS tasks such as ranking a set of documents retrieved for the same query in Information Retrieval (IR) Manning et al. (2008), comparing two answers to the same question in Question Answering (QA) Risch et al. (2021), or measuring the strength of a semantic relation between two entities in Knowledge Graph Completion (KGC) Yoo et al. (2024); Lin et al. (2024).
Figure 1: The two conditions focus on different information described in the two sentences. Human annotators rate the two sentences 1–5, indicating a high-level (5) of semantic textual similarity under $c_{\rm high}$ than $c_{\rm low}$ (1). Our proposed condition-aware sentence embedding (CASE) method reports similarity scores that are well-aligned with the human similarity ratings.
Figure 2: Overview of CASE. An LLM is prompted with $I(s)$ to obtain two separate embeddings $f(c;I(s_{1}))$ and $f(c;I(s_{2}))$ for the same condition $c$ for the two sentences $s_{1}$ and $s_{2}$. The unconditional embedding $f(c;I(\emptyset))$ is then computed using the prompt $I(\emptyset)$ and subtracted from each of those conditional embeddings. Finally, the embeddings are projected to a lower-dimensional space using a supervised FFN and their cosine similarity is computed as the C-STS rating.
We propose CASE (Condition-Aware Sentence Embeddings), an efficient two-step method to combine LLM-based sentence encoders and a lightweight supervised projection, to create the embeddings required for the C-STS task (see Figure 2). In the first step, we prompt the LLM encoders to create the initial embeddings, with an instruction based on conditional semantic similarity to encode the condition given a sentence. Here, the sentence is not encoded explicitly but instead influences the attention scores during token pooling. Compared to Masked Language Model (MLM)-based encoders that have been used in prior work on C-STS, LLM-based encoders can be used as accurate embedding models by direct prompting Tao et al. (2025). They benefit from large-scale datasets and complex architectures with billions of parameters, excelling for various downstream tasks, as demonstrated by the performance on the MTEB leaderboard Muennighoff et al. (2023). However, the best way to use LLMs for C-STS remains elusive, as reported by Lin et al. (2024), who showed that decoder-only LLMs often underperform MLM-based embeddings in C-STS benchmarks. Interestingly, we find that our prompt structure “retrieve [condition], given [sentence]” counter-intuitively performs better than “retrieve [sentence], given [condition]”. Furthermore, a post-processing step of subtracting the embedding of the condition significantly improves the performance and isotropy of the embedding space (as shown in § 4).
Previous C-STS methods fine-tune sentence encoders Li et al. (2024); Tu et al. (2024); Yoo et al. (2024); Lin et al. (2024), which is computationally expensive, particularly for LLM encoders due to their significantly larger number of parameters (>1B), even with parameter-efficient fine-tuning methods. Moreover, LLM encoders produce high-dimensional embeddings (e.g. 4096 for NV-embed-v2 Lee et al. (2024)) compared to MLM encoders (e.g. 768 for SimCSE Gao et al. (2021)), posing challenges when storing and computing inner-products between embeddings. To address these issues, the second step of CASE trains a lightweight supervised Feed Forward Network (FFN) (<5M) in a bi-encoder setting to fine-tune the embeddings in a lower-dimensional space. Qualitative analysis shows that CASE increases the similarity between a sentence and the information emphasised by a given condition, while decreasing the same for information irrelevant to the condition, as expected.
Our CASE method makes the following contributions:
  • It uses the superior semantic capabilities of LLM-based embedding models.
  • It avoids fine-tuning or re-training existing encoders by using a lightweight FFN.
[…]
§ 2 Related Work
Deshpande et al. (2023) proposed the C-STS task and created a dataset containing 18,908 instances where the semantic similarity between two sentences $s_{1}$ and $s_{2}$ is rated under two conditions $c_{\rm high}$ and $c_{\rm low}$ resulting in, respectively, high vs. low similarity between the two sentences. Moreover, they proposed cross-, bi- and tri-encoder baselines. Cross-encoders consider interactions between all tokens in $s_{1}$, $s_{2}$ and $c$, which are computationally expensive for large-scale comparison due to a lack of pre-computed conditional embeddings. Tri-encoders separately encode $s_{1}$, $s_{2}$ and $c$, and then apply late interactions between the condition’s and each sentence’s embeddings, which are also complex and report suboptimal performance. Bi-encoders overcome those limitations by creating a condition-aware embedding for each sentence and computing C-STS efficiently as their inner-product.
Tu et al. (2024) found multiple issues in the dataset created by Deshpande et al. (2023), such as poorly defined conditions and inconsistent similarity ratings for more than half of the dataset. To address this, Tu et al. (2024) re-annotated the validation split from the original C-STS dataset. Moreover, they proposed a QA-based approach for measuring C-STS by first converting conditions into questions, and then using GPT-3.5 to extract the corresponding answers, which are the compared using their embeddings. This QA formulation depends on multiple decoupled components, such as converting conditions into questions, requiring a decoder LLM to extract answers, and using a separate encoder to generate embeddings, which increases the possibility of propagation of errors between components.
Zhang et al. (2025) further investigated the dataset created by Deshpande et al. (2023) and found issues such as problematic condition statements and inaccurate human annotations. They proposed an LLM-based two-step method to improve the dataset by first refining the condition statements and then re-annotating the ratings, creating a larger scale (14176 instances) dataset of better quality, which we use in our experiments as training data.
Alternative architectures to the tri-encoder and cross-encoder have been explored to capture conditional semantics. Yoo et al. (2024) proposed Hyper-CL, a tri-encoder using contrastively learnt hypernetwork Ha et al. (2017) to selectively project the embeddings of $s_{1}$ and $s_{2}$ according to $c$. Hypernetworks introduce an external parameter set that is three times larger than that of the SimCSE model used to encode each sentence, requiring in an excessively large memory space, which is problematic when processing large sets of sentences. Lin et al. (2024) proposed a tri-encoder-based C-STS method where they used routers and heavy-light attention Ainslie et al. (2023) to select the relevant tokens to a given condition. Liu et al. (2025) proposed a conditional contrastive learning method for C-STS, introducing a weighted contrastive loss with a sample augmentation strategy. Li et al. (2024) proposed a cross-encoder approach, which predicts C-STS scores, without creating conditional embeddings. They used a token re-weighting strategy by computing two cross-attention matrices between $(s_{1},s_{2})$ and $c$, which are subsequently used to compute the correlations for the sentence or condition tokens. Although the aforementioned methods improve the performance of cross-encoder and tri -encoder-based C-STS measurement, they still underperform bi-encoders.
Li and Li (2024) proposed BeLLM, which uses backward dependency to enhance LLMs by learning embeddings from uni- to bi-directional attention layers. They fine-tuned the LLaMA2-7B model with their method on the C-STS dataset, outperforming previously proposed fine-tuned models. In particular, they designed a prompt template for this task: “Given the context [condition], summarise the sentence [sentence] in one word:”. They then extracted embeddings from the hidden states of the generated output and used these embeddings to measure sentence similarity. Their results showed that larger model sizes and backward dependencies contribute to the fine-tuning effects. In contrast to our CASE, which uses LLM encoders, they use decoder-only LLMs. Moreover, we do not require token generation and therefore are not restricted to single-word summarisation. Additionally, we find that the reversed structure (“given [sentence], retrieve [condition]”) performs better.
Yamada and Zhang (2025) proposed an unsupervised conditional text embedding method that uses a causal LLM. Specifically, they use the input prompt “Express this text [sentence] in one word in terms of [condition],” and compute the cosine similarity of the last token’s hidden state as the C-STS score. They further validated that instruction-tuned models outperform their non-instruction-tuned counterparts. Similarly, our CASE leverages prompt-based LLM encoders to generate conditional embeddings, yet achieves performance comparable to the best supervised bi-encoder settings in a low-dimensional space.
§ 3 Condition-Aware Sentence Embeddings
An overview of our proposed method is shown in Figure 2, which consists of two-steps. In the first step (§ 3.1), we create two separate embeddings for the condition considering each of the two sentences, one at a time, in the instruction prompt shown to an LLM. Note that it is the condition that is being encoded and the tokens in the sentence (similar to all other tokens in the instruction) are simply modifying the attention scores computed for the tokens in the condition. Intuitively, it can be seen as each sentence filling some missing information required by the condition. However, note that the embeddings obtained from LLMs are not necessarily aligned with the C-STS task. Therefore, in the second step (§ 3.2), we learn a projection layer using the training split from the C-STS dataset. Finally, the C-STS between two sentences is computed as the cosine similarity between the corresponding projected embeddings.
§ 3.1 Extracting Embeddings from LLMs
Given an LLM-based encoder, $f$, we create a $d$-dimensional embedding that captures the semantic relationship between a sentence $s$ and a condition $c$. There are two ways to formulate the input:
  1. 1.
    Encode the condition given the sentence, denoted by $f(c;I(s))$. Here, $I$ is an instruction template that takes $c$ as an argument. We use the following prompt template as I(s)I(s)Retrieve semantically similar texts to a given Condition, given the Sentence : [s], where we substitute $s$ in the placeholder [s]. Next, we provide $c$ as the input text to be encoded by the LLM following the instruction $I(s)$. Finally, the token embeddings of $c$ are aggregated according to one of the pooling methods to create $f(c;I(s))$ (different pooling methods are discussed in Appendix D).
  2. 2.
    Encode the sentence given the condition, denoted by $f(s;I(c))$. Here we swap the sentence and the condition in the above formulation, recalling that both $s$ and $c$ are text strings. In particular, as shown later in our experiments (§ 4.2), comparing the embedding for $c$ created given $s_{1}$ and $s_{2}$ results in better performance on the C-STS benchmark for all LLM encoders. For this reason, we keep the best-performing $f(c;I(s))$ in Figure 2 and our method for subsequent experiments. This is because $s_{1}$ and $s_{2}$ will also contain information irrelevant to $c$, which will affect the cosine similarity computed between $f(s_{1};I(c))$ and $f(s_{2};I(c))$. On the other hand, the cosine similarity between $f(c;I(s_{1}))$ and $f(c;I(s_{2}))$ is a more accurate estimate of C-STS between $s_{1}$ and $s_{2}$ under $c$ because it is purely based on contextualised representation shifts of $c$ conditioned on $s_{1}$ and $s_{2}$ separately.
Condition statements contain their own intrinsic semantics (e.g., the phrase colour of dress), which is irrelevant when distinguishing specific semantic values based on sentences (e.g., orange vs. grey). To address this problem, we introduce a post-processing step of subtracting the embedding of the condition to reduce the effect of tokens in the condition that are irrelevant to the sentence, thereby consistently improving the accuracy of the condition-aware embeddings. Specifically, we use the prompt $I(\emptyset)$ Retrieve semantically similar texts to obtain the embedding of a condition $c$, and denote this by $f(c;I(\emptyset))$. As we see later in our experiments, by subtracting $f(c;I(\emptyset))$ from $f(c;I(s))$, we improve the performance on the C-STS task and the isotropy of embeddings. This first step is fully unsupervised and a zero-shot prompt template is used as $I$.
§ 3.2 Supervised Projection Learning
The LLM embeddings computed in § 3.1 has two main drawbacks. First, although LLMs are typically trained on massive text collections and instruction-tuned for diverse tasks Muennighoff et al. (2023), their performance on C-STS tasks have been poor Lin et al. (2024). As seen from our condition-aware prompt template, an LLM must be able to separately handle a variable condition statement and a fixed instruction. This setup is different from most tasks on which LLMs are typically trained on, where the instruction remains fixed across all inputs. Second, relative to MLMs-based sentence embeddings, LLMs produce much higher dimensional embeddings, which can be problematic due to their memory requirements (especially when operating on a limited GPU memory) and the computational cost involved in inner-product computations. In tasks such as dense retrieval, we must compare millions of documents against a query to find the nearest neighbours under strict latency requirements, and low-dimensional embeddings are preferable.
To address the above-mentioned drawbacks, we propose a supervised projection learning method. Specifically, we freeze the model parameters of the LLM and use a FFN layer that takes $f(c;I(s))$ as the input and returns a $k$-dimensional ($k\leq d)$ embedding $g(f(c;I(s);\theta)$, where $\theta$ denotes the learnable parameters of the FFN. Finally, we define ${\rm CASE}(s,c)$ as the projection of the offset between the conditional and the unconditional embeddings of $c$ under $s$, given by,
${\rm CASE}(s,c)=g(f(c;I(s))-f(c;I(\emptyset));\theta)$ (1)
We use the human-annotated similarity ratings $r$ in the C-STS train instances $\mathcal{D}$ to learn $\theta$. Specifically, we minimise the squared error between the human ratings and the cosine similarity computed using the corresponding CASE as given by (2).
$\sum_{(s_{1},s_{2},c,r)\in\mathcal{D}}\left(\cos({\rm CASE}(s_{1},c),{\rm CASE}(s_{2},c))-r\right)^{2}$ (2)
Here, $\cos$ denotes the cosine similarity between the projected embeddings. We use Adam optimiser Kingma and Ba (2014) to find the optimal $\theta$ that minimises the loss given by (2). Recall that only the FFN parameters are updated during this projection learning step, while keeping the parameters of the LLM fixed, which makes it extremely efficient. Using the learnt projection, we compute the C-STS between $s_{1}$ and $s_{2}$ under $c$ as the cosine similarity between the embeddings ${\rm CASE}(s_{1},c)$ and ${\rm CASE}(s_{2},c)$.
§ 4 Experiments and Results
To evaluate the effectiveness of the LLM-based and MLM-based sentence embeddings as described in § 3, we use six sentence encoders, out of which three are LLM-based: NV-Embed-v2 (4096 dimensional NV), SFR-Embedding-Mistral (4096 dimensional SFR), gte-Qwen2-7B-instruct (3584 dimensional GTE) and three are MLM-based: Multilingual-E5-large-instruct (1024 dimensional E5), sup-simcse-roberta-large (1024 dimensional SimCSE_large), and sup-simcse-bert-base-uncased (768 dimensional SimCSE_base). Further details provided in Appendix A.
Model sent/cond? Spear.
NV sent - c 16.98
sent 22.07
cond - c 37.61
cond 33.85
SFR sent - c 19.54
sent 11.89
cond - c 20.44
cond 18.32
GTE sent - c 7.16
sent 7.16
cond - c 20.40
cond 16.58
E5 sent - c 11.08
sent 3.77
cond - c 15.37
cond 6.07
SimCSE_large $\operatorname{CONC}(s+c)$ 5.59
$\operatorname{CONC}(c+s)$ 4.00
$\operatorname{CONC}(s+c)$ - c 8.32
$\operatorname{\textbf{CONC}}\bm{(c+s)}$ - c 8.58
SimCSE_base $\operatorname{CONC}(s+c)$ 4.37
$\operatorname{CONC}(c+s)$ 1.25
$\operatorname{\textbf{CONC}}\bm{(s+c)}$ - c 7.05
$\operatorname{CONC}(c+s)$ - c 6.00
Table 1: Spearman scores for different sentence embedding models and encoding settings.
We evaluate model performance on different pooling methods, prompt settings, and sentence constructions. Moreover, we evaluate the supervised methods of linear and non-linear FFNs. Both the linear and non-linear FFNs are Siamese bi-encoders with weight-sharing. As explained in § 3.2, they take $f(c;I(s_{1}))$ and $f(c;I(s_{2}))$ as the input embeddings, and return the projected embeddings ${\rm CASE}(s_{1},c)$ and ${\rm CASE}(s_{2},c)$ as the outputs, which can be compared using a similarity metric such as the cosine similarity.
Linear FFN performs a linear transformation:
$\mathbf{z}=\text{\bf W}\bm{e}$ (3)
where $\text{\bf W}\in\mathbb{R}^{d^{\prime}\times d}$ is the learned projection matrix.
Our non-linear FFN is a one-layer FFN with a activation function LeakyReLU:
$\mathbf{z^{\prime}}$ $=\operatorname{Dropout}\left(\operatorname{LeakyReLU}\left(\text{\bf W}^{\prime}\bm{e}\right)\right)$ (4)
A dropout layer is applied to reduce any overfitting Hinton et al. (2012). Hyperparameters are tuned on a held-out validation set. We select a learning rate of $10^{-3}$ and a batch size of 512. The dropout rate is set to $15\%$ for non-linear FFN. Appendix B provides the ablation study for the architectures of FFNs.
To address annotation errors in the original C-STS dataset such as ambiguous and invalid conditions, Tu et al. (2024) re-annotated the original validation set. Zhang et al. (2025) further identified the issues on both condition statements and ratings. They improved the dataset by first revising the condition statements and then re-annotating the similarity ratings for both the training and validation sets. To conduct a more accurate and reliable evaluation, we use the most recent revised and cleaned datasets by Zhang et al. (2025). [^2] https://huggingface.co/datasets/LivNLP/C-STS-Reannotated We split their validation set 70% vs. 30% randomly to a validation set of 1983 instances and a test set of 851 instances. In this C-STS dataset, there are 14176 unique instances with 4383 unique conditions. As for each sentence pair ($s_{1}$, $s_{2}$), there are two conditions ($c_{\rm high}$ and $c_{\rm low}$). There are 7088 unique sentence pairs, of which 5719 pairs have unique condition pairs ($c_{\rm high}$, $c_{\rm low}$). In other words, 19.3% of the sentence pairs share their conditions with other sentence pairs.
We use a single p3.24xl EC2 instance (8x V100 GPUs) for learning sentence embeddings, and a separate NVIDIA RTX A6000 GPU for supervised projection learning. Pytorch 2.0.1 with cuda 11.7 is used for FFN projection. These settings are fixed across all experiments. For reducing 4096-dimensional LLM-based sentence embeddings to 512-dimensional, training a non-linear FFN for CASE takes less than 1 minute (wall-clock time).
§ 4.1 Evaluation Metrics
We evaluate the performance of a sentence encoder by the Spearman correlation coefficient between the cosine similarity scores computed using the embeddings produced by an encoder and the human similarity ratings on the test set.
§ 4.2 C-STS Measurement
Model $I_{\text{iso}}$ (-c) $I_{\text{iso}}$
NV 0.9611 0.9471
SFR 0.9463 0.9266
GTE 0.9490 0.9310
E5 0.8906 0.8368
SimCSE_base 0.9461 0.9297
SimCSE_large 0.9499 0.9406
Table 2: $I_{\text{iso}}$ values for each model. The left column of $I_{\text{iso}}$ (-c) lists values with the post-processing step of subtracting the embedding of condition c. $I_{\text{iso}}$ values close to 1 indicate high isotropy.
To generate CASE, we apply different ways to construct the prompt for LLM-based embeddings and to concatenate the condition and sentence for MLM-based embeddings. For LLM embeddings, we have two main settings: (a) $\operatorname{cond}=f(c;I(s))$, where we encode the condition given the sentence, and (b) $\operatorname{sent}=f(s;I(c))$, where we encode the sentence given the condition as explained in § 3.1. For MLM embeddings, we evaluate the two settings: (a) $\operatorname{CONC}(c+s)$, where we concatenate sentence after the condition, and (b) $\operatorname{CONC}(s+c)$, where we concatenate condition after the sentence. For each setting, we evaluate the effect of subtracting the condition embedding, $\bm{c}=f(c;I(\emptyset))$.
(a) Test Spearman for linear FFN
(b) Test Spearman for non-linear FFN
Figure 3: Spearman correlation on test set for different models over training steps with dimensionality 512. The $y$-axes of both subfigures are aligned, facilitating a direct comparison of the Spearman coefficients across the two line charts, with the same colour for the same model. Best viewed in colour.
The test performance for different settings and models are shown in Table 1. For LLM embeddings, cond consistently reports higher Spearman correlation than sent, suggesting that embedding the condition given the sentence is more effective for C-STS measurement than embedding the sentence given the condition. Moreover, we see that subtracting $\bm{c}$ further improves performance across all settings. The former approach reduces the noise due to the tokens in a sentence, which are irrelevant to the given condition. For MLM-based embeddings, subtracting $\bm{c}$ also improves performance. This operation removes the sentence-independent components of the condition, thereby isolating the contrastive semantics — the aspects of the condition that are actually altered by the sentence. As a result, CASE focuses on the semantic variation between the two sentences under the given condition, rather than the redundant information shared across all condition phrases.
Additionally, we discovered that subtracting $\bm{c}$ in a post-processing step improves isotropy of the embeddings (full details in Appendix C). We use the approximated isotropy metric $I_{\text{iso}}$ Durdy et al. (2023) as a numerically stable method to measure isotropy. $I_{\text{iso}}$ is estimated by randomly sampling unit vectors on the hypersphere and computing the ratio of their minimum to maximum alignment with the embeddings across random directions. Table 2 shows that the post-processing step of subtracting the condition $\bm{c}$ gives higher $I_{\text{iso}}$ values, indicating the improvement of isotropy. This makes subtle semantic differences between sentences under a given condition more distinguishable. This is in-line with prior work reporting a positive correlation between isotropy and improved performance in embedding models Rajaee and Pilehvar (2022); Su et al. (2021). The effect of pooling method on performance is discussed in Appendix D, where we find the latent pooling in NV to perform best. Therefore, we use the ($\operatorname{cond}-\bm{c}$) setting (which reports the best performance) for the six sentence encoders to conduct the subsequent experiments.
Model Non-linear FFN Linear FFN
NV 74.94 73.55
SFR 68.36 59.02
GTE 64.16 55.88
E5 65.07 48.25
SimCSE_large 63.64 48.15
SimCSE_base 61.52 42.22
Table 3: Spearman correlation of embedding models with supervised methods at reduced dimensionality $512$.
Figure 4: Spearman correlation coefficients of our CASE for NV embeddings on the test set over different dimensionalities. The horizontal dashed line represents the performance of fine-tuned SimCSE-base (768-dimensional).
We show the training curves for our supervised FFNs in Figure 3. Overall, the non-linear FFNs achieve significantly higher Spearman correlation than the linear FFNs. Moreover, non-linear FFNs converge faster, typically reaching their peak performance within 20 epochs. The performance of the linear FFNs, gradually increases and eventually converges as the training progresses. NV consistently performs the best for both linear and non-linear FFNs.
We compare sentence encoders in Table 3 for a fixed (i.e. 512) dimensional projection from their original embeddings. NV obtains the highest Spearman coefficient across all sentence encoders and supervised methods. GTE and SFR also achieve high Spearman coefficients under the non-linear FFNs. SimCSE_large and SimCSE_base have the lowest Spearman coefficients. Overall, we see that LLM-based embeddings perform better than the MLM-based embeddings on the C-STS task. Importantly, all embeddings are projected to the same lower dimensionality.
Method Spearman dim #params
Deshpande et al.
w/ SimCSE-large 79.51 1024 354.3M
w/ SimCSE-base 72.12 768 124.1M
CASE
w/ non-linear FFN 74.94 512 2.1M
Table 4: Comparison of Spearman correlation, embedding dimensionality, and trainable parameters of previously proposed methods and our CASE. #params denotes the number of trainable parameters.
s1: Young woman in orange dress about to serve in tennis game, on blue court with green sides.
s2: A girl playing tennis wears a gray uniform and holds her black racket behind her.
[…]
s1: Two snow skiers with ski poles and snow skis, standing on top of a mountain with other skiers around them.
s2: A skier stands alone at the top of a snowy slope with blue skies and mountains in the distance.
[…]
s1: A bunch of people standing around at the beach with a kite in the air.
s2: a beach scene with a beach chair decorated with the Canadian Flag and surfers walking by with their surfboards
[…]
Table 5: Example of similarity scores for two conditions applied to the same sentence pair, based on non-linear FFN with dimensionality $512$ on NV. The table shows how supervised FFN improves the CASE for C-STS task. cos(,)($\cdot$$\cdot$) denotes cosine similarity. Answer 1 and Answer 2 refer to the corresponding answers for Sentence 1 and Sentence 2 under conditions. The predicted similarity scores before and after applying supervised FFN are listed on the left and right of the arrow. That is, the similarity score for original high-dimensional LLM-based embeddings is on the left of the arrow, while the similarity score for CASE is on the right.
To study how the performance of the proposed CASE varies with the dimensionality of the supervised projection, we plot the Spearman correlation measured on the test split of the C-STS dataset in Figure 4 using NV embeddings. Because we are specifically interested in examining how CASE performs in the low-dimensional regime, which is attractive from a computational viewpoint, we consider dimensionalities in range [50, 1000]. Moreover, as a reference, we plot the performance of the previously proposed bi-encoder model by Deshpande et al. (2023) for the SimCSE-base model, which is the smallest dimensional (768) encoder used in their work. From Figure 4 we see that for dimensionalities greater than 200, CASE consistently outperforms SimCSE-base bi-encoder. This result has practical implications because it shows that we can reduce the high dimensionality of LLMs-based (NV) embeddings from 4096 to 512 with an 8X compression, while preserving its high performance. This result is consistent with Matryoshka Representation Learning (MRL) Kusupati et al. (2022), where they show that small-dimensional embeddings can maintain high performance in downstream tasks. MRL is applicable with our MLM and LLM embeddings to further reduce the dimensionality, while our FFNs project embeddings into a specific lower dimensionality rather than the nested granularities of MRL. Our results show that embeddings can be efficiently compressed in the bi-encoder setting on the C-STS task.
We compare CASE against the best bi-encoder models with fine-tuned SimCSE by Deshpande et al. (2023). As shown in Table 4, our proposed CASE with non-linear FFN achieves 74.9% Spearman correlation, outperforming fine-tuned SimCSE_base while using only 2.1M trainable parameters and a 512-dimensional embedding space. Compared with SimCSE_large, CASE maintains 94% of its performance, but requires 50% fewer dimensions and only 0.6% of its trainable parameters. This highlights that CASE is a simple but efficient method.
§ 4.3 Case Study
To further illustrate how CASE captures condition-dependent semantics, Table 5 presents qualitative examples. Similarity scores are computed and compared for three sentence pairs for the original high-dimensional LLM-based embeddings and low-dimensional trained CASE. By comparing the similarity scores with and without the supervised projection, we can evaluate the ability of CASE to focus on the condition-related information. Compared to the unconditional similarity between the two sentences, conditional similarities scores computed using NV embeddings align well with the human ratings. For example, in the top row in Table 5, we see that the unconditional similarity between the two sentences reduces from $0.5006$ to $0.4747$ under $c_{1}$, while increasing to $0.6233$ under $c_{2}$. Moreover, the conditional similarities are further appropriately amplified by CASE using the supervised projection (i.e. decreasing to $0.2522$ under $c_{1}$, while increasing to $0.9660$ under $c_{2}$). Specifically, CASE reduces the similarity between the two sentences under the lower-rated condition, while increasing the same under the high-rated condition.
We also treat the conditions as questions for sentences and extract the relevant information as answers to compare whether CASE can focus on the condition-related information. The overall similarity trend is consistent with the actual ratings. For all sentence and answer pairs, the similarity scores increase after supervised projection. This demonstrates that CASE can effectively capture the shift in conditional meaning between sentences under different conditions.
Method Spearman Dim
Qwen3-8B (frozen)
     w/o FFN 24.59 4096
     w/ FFN (CASE) 71.64 512
Qwen3-8B (LoRA fine-tuned)
     w/o FFN 79.91 4096
     w/ FFN 79.11 512
Table 6: Comparison of fine-tuning. We report the Spearman correlation of the original and LoRA fine-tuned Qwen3-8B encoders, both with and without our supervised non-linear FFN.
§ 4.4 Comparision with Fine-tuning LLM Encoders
We investigate the effects of fine-tuning LLM encoders using LoRA Hu et al. (2022), and compare its performance with our CASE method. Specifically, we use the high-performing encoder Qwen3-Embedding-8B (Qwen3-8B) [^3] https://huggingface.co/Qwen/Qwen3-Embedding-8B in the MTEB learderboard. The encoder uses a 4096-dimensional embedding space with last-token pooling. We fine-tune the Qwen3-8B on the C-STS dataset with Pearson correlation loss computed over the cosine similarities. For the LoRA configuration, we use a rank $r=32$, a scaling factor $\alpha=32$, and a dropout rate of $0.05$. Adaptation is applied to the query, key, value, and output projection layers of the attention mechanism. Fine-tuning Qwen3-8B requires over four hours of wall-clock time on two NVIDIA A100 (80GB) GPUs. In contrast, CASE learns a supervised FFN on pre-computed embeddings and completes training in under one minute on a single NVIDIA RTX A6000 GPU (approximately 10 seconds on an A100). Moreover, LoRA fine-tuning of 8B-scale models requires substantial memory for model loading and activation storage, often necessitating at least 24 GB of GPU memory in practical settings, while CASE supervision requires negligible GPU memory usage (<1 GB).
[…]
§ 5 Conclusion
We propose CASE, a method for computing the semantic textual similarity between two sentences under a given condition. CASE encodes the condition under a sentence, and then subtracts the embedding of the condition in a post-processing step. Our experimental results show that LLM embeddings consistently outperform MLM embeddings for C-STS. Moreover, we introduce an efficient supervised projection learning method to improve the performance in C-STS while projecting embeddings to lower-dimensional spaces for efficient similarity computations.
Acknowledgements
Danushka Bollegala holds concurrent appointments as a Professor at the University of Liverpool and as an Amazon Scholar at Amazon. This paper describes work performed at the University of Liverpool and is not associated with Amazon.
§ 6 Limitations
Our evaluations cover only English, which is a morphologically limited language. To the best of our knowledge, C-STS datasets have not been annotated for languages other than English, which has forced all prior work on C-STS to conduct experiments using only English data. However, the sentence encoders we used in our experiments support multiple languages. Therefore, we consider it to be an important future research direction to create multilingual datasets for C-STS and evaluate the effectiveness of our proposed method in multilingual settings.
There is a large number of sentence encoders (over 18,000 sentence encoders as in January 2026 evaluated in Hugging Face Hub [^4] https://huggingface.co/models). However, due to computational limitations, we had to select a subset covering the best performing (top-ranked on MTEB leaderboard at the time of writing) models for our evaluations.
§ 7 Ethical Concerns
We did not collect or annotate any datasets in this project. Instead, we use existing C-STS datasets annotated and made available by Deshpande et al. (2023), Tu et al. (2024), and Zhang et al. (2025). To the best of our knowledge, no ethical issues have been raised regarding those datasets.
We use multiple pre-trained and publicly available MLM- and LLM-based sentence encoders Kaneko and Bollegala (2021). Both MLMs and LLMs are known to encode unfair social biases such as gender or racial biases. We have not evaluated how such social biases would be influenced by the CASE learning method proposed in this work. Therefore, we consider it would be important to measure the social biases in CASE created in this work before they are deployed in real-world applications.
Supplementary Materials
Appendix A Models
To evaluate the effectiveness of the LLM-based and MLM-based sentence embeddings as described in § 3, we apply six sentence encoders, out of which three are LLM-based: NV-Embed-v2 (4096 dimensional and uses latent pooling) (NV) [^5] https://huggingface.co/nvidia/NV-Embed-v2, SFR-Embedding-Mistral (4096 dimensional and uses average pooling) (SFR) [^6] https://huggingface.co/Salesforce/SFR-Embedding-Mistral, gte-Qwen2-7B-instruct (3584 dimensional and uses last token pooling) (GTE) [^7] https://huggingface.co/Alibaba-NLP/gte-Qwen2-7B-instruct and three are MLM-based: Multilingual-E5-large-instruct (1024 dimensional and uses average pooling) (E5) [^8] https://huggingface.co/intfloat/multilingual-e5-large-instruct, sup-simcse-roberta-large (1024 dimensional and uses last token pooling) (SimCSE_large) [^9] https://huggingface.co/princeton-nlp/sup-simcse-roberta-large, and sup-simcse-bert-base-uncased (786 dimensional and uses last token pooling) (SimCSE_base) [^10] https://huggingface.co/princeton-nlp/sup-simcse-bert-base-uncased.
Appendix B Ablation Study for Supervised FFN
Table 7 provides the ablation study of linear and non-linear FFN, using NV as a representative model (for other encoders, we observe similar trends). We assess the impact of including a dropout layer. For non-linear FFN, we further compare the effects of different activation functions (ReLU, GeLU, SiLU, LeakyReLU) and the number of layers (one, two). Note that based on preliminary experiments, we set the dropout rate to 20% for the Linear FFN and 15% for the Non-linear FFNs as the best configuration. Additionally, for two-layer non-linear FFNs, the intermediate layer dimension is set to 1024, which performs best among the candidate set {256, 768, 1024, 2048, 8192}.
Linear FFN without dropout gives higher performance than that with dropout, with a Spearman correlation of 73.55. For non-linear FFNs, one-layer FFNs generally outperform two-layer FFNs by 1-3 points. Additionally, introducing a dropout layer improves the performance, probably by reducing the impact of overfitting. Among the four activation functions, LeakyReLU yields the best results for both two-layer and one-layer FFNs, achieving a Spearman correlation over 74. Consequently, we proceed with the linear FFN (without dropout) and the non-linear FFN (one layer, activation function LeakyReLU, dropout of 15%) for the subsequent experiments.
Activation w/ Dropout w/o Dropout
[…]
72.90 73.55
[…]
   ReLU 73.93 71.47
   GeLU 67.66 69.06
   SiLU 69.33 69.58
   LeakyReLU 74.05 73.29
[…]
   ReLU 73.27 71.56
   GeLU 71.05 70.56
   SiLU 71.08 71.13
   LeakyReLU 74.94 71.16
Table 7: Ablation study comparing Linear and Non-linear FFNs. Numbers in parentheses indicate layer dimensions.
Appendix C Isotropy of Embeddings
[…]
From Table 8, Figure 5, and Figure 6, we see that the embeddings after subtracting $\bm{c}$ have a lower mean cosine similarity to the mean vector and a higher standard deviation, indicating that they are more spread out in the embedding space. In contrast, the embeddings without subtracting $\bm{c}$ are more clustered around a central direction (higher mean, lower standard deviation), reflecting anisotropy, a tendency for vectors to concentrate in a narrow region. Therefore, embeddings after subtracting c tend to be more isotropic, indicating better distributional diversity.
Model Embedding Type Mean Std
NV cond - c 0.407 0.084
cond 0.492 0.067
SFR cond - c 0.537 0.055
cond 0.708 0.021
GTE cond - c 0.489 0.067
cond 0.696 0.036
E5 cond - c 0.542 0.056
cond 0.897 0.010
SimCSE_base $\operatorname{CONC}(c+s)$ - c 0.254 0.095
$\operatorname{CONC}(c+s)$ 0.347 0.088
SimCSE_large $\operatorname{CONC}(c+s)$ - c 0.248 0.110
$\operatorname{CONC}(c+s)$ 0.333 0.085
Table 8: Cosine similarity to mean vector: comparing mean and standard deviation of two embedding types across three LLM-based and three MLM-based models.
We use an efficient approximation method to compute the isotropy Durdy et al. (2023). Given a normalized embedding matrix $E\in\mathbb{R}^{n\times d}$ where each row represents a unit vector, instead of computing the eigenvectors of the covariance matrix, we randomly sample $k$ unit vectors $\bm{u}_{1},\bm{u}_{2},...,\bm{u}_{k}$ from the unit hypersphere $\mathbb{S}^{d-1}$. For each sampled direction $\bm{u}_{i}$, we compute the function $F(\bm{u}_{i})=\sum_{j=1}^{n}\exp(\bm{e}_{j}{}^{\top}\bm{u}_{i})$, where $\bm{e}_{j}$ represents the $j$-th embedding vector. The estimated isotropy is then defined as the ratio between the minimum and maximum values of $F$, given by (5).
$I_{\text{iso}}\approx\frac{\min_{i\in{1,...,k}}F(u_{i})}{\max_{i\in{1,...,k}}F(u_{i})}$ (5)
$I_{\text{iso}}$ values close to 1 indicates that the embedding space shows high isotropy where vectors are uniformly distributed in all directions in the high-dimensional space instead of clustering in some small number of dominant directions. In contrast, when $I_{\text{iso}}$ approaches 0, it means a significant anisotropy in the embedding space. We use $k=1000$ random directions to compute the approximation of isotropy. Table 9 shows that the post-processing step of subtracting the condition c gives higher $I_{\text{iso}}$ values, indicating the improvement of isotropy. Intuitively, when the embedding space is isotropic it becomes easier to differentiate between smaller similarity differences, thus improving the C-STS estimation.
Model $I_{\text{iso}}$ (-c) $I_{\text{iso}}$
NV 0.9611 0.9471
SFR 0.9463 0.9266
GTE 0.9490 0.9310
E5 0.8906 0.8368
SimCSE_base 0.9461 0.9297
SimCSE_large 0.9499 0.9406
Table 9: Isotropy values $I_{\text{iso}}$ for each model. The left column of $I_{\text{iso}}$ (-c) lists values with post-processing step of subtracting the condition c.
Model Pooling sent/cond? Spear.
NV latent sent - c 16.98
latent sent 22.07
latent cond - c 37.61
latent cond 33.85
SFR last sent - c 11.88
last sent -0.18
last cond - c 19.28
last cond 13.00
average sent - c 19.54
average sent 11.89
average cond - c 20.44
average cond 18.32
GTE last sent - c 7.16
last sent 7.16
last cond - c 20.40
last cond 16.58
average sent - c 13.01
average sent 11.34
average cond - c 17.87
average cond 18.42
E5 average sent - c 11.08
average sent 3.77
average cond - c 15.37
average cond 6.07
last sent - c 9.28
last sent 0.49
last cond - c 8.90
last cond 3.05
SimCSE_large $\operatorname{CONC}(s+c)$ 5.59
$\operatorname{CONC}(c+s)$ 4.00
$\operatorname{CONC}(s+c)$ - c 8.32
$\operatorname{\textbf{CONC}}\bm{(c+s)}$ - c 8.58
SimCSE_base $\operatorname{CONC}(s+c)$ 4.37
$\operatorname{CONC}(c+s)$ 1.25
$\operatorname{\textbf{CONC}}\bm{(s+c)}$ - c 7.05
$\operatorname{CONC}(c+s)$ - c 6.00
Table 10: Spearman scores for sentence embedding models. The original dimensionality of each model is indicated in parentheses. latent denotes latent attention pooling for NV, whereas last and average correspond to last token pooling and average pooling, respectively.
Appendix D Full Results for all Models
Each LLM encoder has its own optimal pooling method, recommended by the original authors of those models. From Table 10 we see that the performance varies significantly depending on the pooling method being used, while the latent attention pooling used in NV reporting the best results. Note that NV does not support last or average pooling, while latent pooling is not supported by the other models.
(a) NV: cond - c
(b) NV: cond
(c) SFR: cond - c
(d) SFR: cond
(e) GTE: cond - c
(f) GTE: cond
Figure 5: Embeddings-to-mean cosine similarity distributions across three LLM-based models. Each row compares cond - c and cond representations.
(a) E5: cond - c
(b) E5: cond
(c) SimCSE_base: $\operatorname{CONC}(c+s)$ - c
(d) SimCSE_base: $\operatorname{CONC}(c+s)$
(e) SimCSE_large: $\operatorname{CONC}(c+s)$ - c
(f) SimCSE_large: $\operatorname{CONC}(c+s)$
Figure 6: Embeddings-to-mean cosine similarity distributions across three MLM-based models. Each row compares two embedding types.